OpenStudy (clongoria2):

Can someone check my work? Derive the equation of the parabola with a focus at (6, 2) and a directrix of y = 1.

OpenStudy (clongoria2):

(x-6)^2 + y^2 - 4y + 4 = y^2 - 2y + 1 (x-6)^2 + y^2 - 4y = y^2 - 2y - 3 (x-6)^2 + y^2 = y^2 + 2y - 3 (x-6)^2 = + 2y - 3 1/2(x-6)^2 = - 3/2 f(x) = 1/2(x-6)^2 + 3/2

OpenStudy (clongoria2):

I think i followed the steps correctly. Its just a little hard to know if i did it right.

Nnesha (nnesha):

do you need it in vertex form ?

OpenStudy (clongoria2):

The answers are A. f(x) = -1/2 (x-6)^2 + 3/2 B. f(x) = 1/2 (x-6)^2 + 3/2 C. f(x) = -1/2 (x+3/2)^2 + 6 D. f(x) = 1/2 (x+3/2)^2 + 6

OpenStudy (clongoria2):

I chose B. I just wanted to know if i did it right.

Nnesha (nnesha):

yes correct good job!

OpenStudy (clongoria2):

Thanks

Nnesha (nnesha):

your answer is correct but i think u skipped one step (x-6)^2 = + 2y - 3 first add (x-6)^2+3=2y now divide by 2 your work ` (x-6)^2 = + 2y - 3` `1/2(x-6)^2 = - 3/2` looks like you divided by 2 first and then move the constant to the left side great work keep it up!